# Information Science Principles of Machine Learning: A Causal Chain Meta-Framework Based on Formalized Information Mapping

### v9

@@ A.1 A Simple Feedforward Neural Network Model
$W_{1}=\begin{bmatrix}w_{11}^{1}&w_{12}^{1}\\ w_{21}^{1}&w_{22}^{1}\\ w_{31}^{1}&w_{32}^{1}\end{bmatrix},\quad b_{1}=\begin{bmatrix}b_{1}^{1}\\ b_{2}^{1}\\ b_{3}^{1}\end{bmatrix}$ (A.3)
$W_{2}=\begin{bmatrix}w_{11}^{2}&w_{12}^{2}&w_{13}^{2}\\ w_{21}^{2}&w_{22}^{2}&w_{23}^{2}\end{bmatrix},\quad b_{2}=\begin{bmatrix}b_{1}^{2}\\ b_{2}^{2}\end{bmatrix}$ (A.4)
@@
$X=\begin{bmatrix}x_{1}\\ x_{2}\end{bmatrix},\quad Y=\begin{bmatrix}y_{1}\\ y_{2}\end{bmatrix},\quad Y_{t}=\begin{bmatrix}y_{t1}\\ y_{t2}\end{bmatrix}\quad and\quad Y_{t}=\text{Target}(X)$ (A.5)
$L=\text{MSE}(Y,Y_{t})=\frac{1}{2}\|Y-Y_{t}\|^{2}=\frac{1}{2}((y_{1}-y_{t1})^{2}+(y_{2}-y_{t2})^{2})$ (A.6)
@@
$\nabla_{W_{2}}L=(Y-Y_{t})(\text{ReLU}(W_{1}X+b_{1}))^{T}$ (A.7)
$\nabla_{b_{2}}L=(Y-Y_{t})$ (A.8)
$\nabla_{W_{1}}L=(W_{2}^{T}(Y-Y_{t})\odot\text{ReLU}^{\prime}(W_{1}X+b_{1}))X^{T}$ (A.9)
$\nabla_{b_{1}}L=W_{2}^{T}(Y-Y_{t})\odot\text{ReLU}^{\prime}(W_{1}X+b_{1})$ (A.10)
In equations (A.9) and A.10, the operator $\odot$ denotes element-wise multiplication, and ReLU’ represents the derivative of the ReLU function. Since ReLU is not differentiable at zero, we additionally define ReLU’(0) = 0. Thus, the update formulas for optimizing the weight matrices and bias vectors during M’s training are:
$W_{1}^{\text{new}}$ $=W_{1}-\eta\nabla_{W_{1}}L,\quad b_{1}^{\text{new}}=b_{1}-\eta\nabla_{b_{1}}L$ (A.11)
$W_{2}^{\text{new}}$ $=W_{2}-\eta\nabla_{W_{2}}L,\quad b_{2}^{\text{new}}=b_{2}-\eta\nabla_{b_{2}}L$ (A.12)
@@
A.2 Symbol Definitions
MLT-MF uses a formal system to express the operational logic of each phase throughout the model’s full lifecycle. Defining the symbols is a prerequisite for formal modeling. To this end, Tables A1A4 provide definitions for the constants, variables, functions, and predicates used in the formal representation of model M.
Table A1: Definition of Constants
@@ A.2 Symbol Definitions
Table A2: Definition of Variables
@@
$\varphi_{o_{i}}$ WFFs well-formed formulas contained in the informational ontological state of M’s initial model
$\psi_{o_{i}}$ WFFs well-formed formulas contained in the informational carrier state of M’s initial model
$\varphi_{t_{i}}$ WFFs well-formed formulas contained in the training informational ontological state
$\psi_{t_{i}}$ WFFs well-formed formulas contained in the training informational carrier state
$\varphi_{u_{i}}$ WFFs well-formed formulas contained in the informational ontological state of M’s optimized model
$\psi_{u_{i}}$ WFFs well-formed formulas contained in the informational carrier state of M’s optimized model
$\varphi_{q_{i}}$ WFFs well-formed formulas contained in the input informational ontological state of M
$\psi_{q_{i}}$ WFFs well-formed formulas contained in the input informational carrier state of M
$\varphi_{r_{i}}$ WFFs well-formed formulas contained in the output informational ontological state of M
$\psi_{r_{i}}$ WFFs well-formed formulas contained in the output informational carrier state of M
$\psi_{v_{i}}$ WFFs well-formed formulas contained in the informational ontological state of M’s re-optimized model
$\psi_{v_{i}}$ WFFs well-formed formulas contained in the informational carrier state of M’s reoptimized model
@@
$W_{k}^{(o)}$ matrix weight matrix from the k-th to the (k+1)-th layer in the initial model
$w_{ij}^{(o)k}$ real number weight from the j-th neuron in the k-th layer to the i-th neuron in the (k+1)-th layer in the initial model
$b_{k}^{(o)}$ vector bias vector of neurons in the k-th layer in the initial model
$b_{i}^{(o)k}$ real number bias of the i-th neuron in the k-th layer in the initial model
$W_{k}^{(u)}$ matrix weight matrix from the k-th to the (k+1)-th layer in the optimized model
$w_{ij}^{(u)k}$ real number weight from the j-th neuron in the k-th layer to the i-th neuron in the (k+1)-th layer in the optimized model
$b_{k}^{(u)}$ vector bias vector of neurons in the k-th layer in the optimized model
$b_{i}^{(u)k}$ real number bias of the i-th neuron in the k-th layer in the optimized model
$W_{k}^{(v)}$ matrix weight matrix from the k-th to the (k+1)-th layer in the re-optimized model
$w_{ij}^{(v)k}$ real number weight from the j-th neuron in the k-th layer to the i-th neuron in the (k+1)-th layer in the re-optimized model
$b_{k}^{(v)}$ vector bias vector of neurons in the k-th layer in the re-optimized model
$b_{i}^{(v)k}$ real number bias of the i-th neuron in the k-th layer in the re-optimized model
$W_{k}^{(w)}$ matrix weight matrix from the k-th to the (k+1)-th layer in the further re-optimized model
$w_{ij}^{(w)k}$ real number weight from the j-th neuron in the k-th layer to the i-th neuron in the (k+1)-th layer in the further re-optimized model
$b_{k}^{(w)}$ vector bias vector of neurons in the k-th layer in the further re-optimized model
$b_{i}^{(w)k}$ real number bias of the i-th neuron in the k-th layer in the further re-optimized model
@@
$x_{i}$ real number the i-th element of the input vector
$h_{i}$ real number the i-th element of the hidden layer vector
$y_{i}$ real number the i-th element of the output vector
$y_{ti}$ real number the i-th element of the output reference vector
@@
$\nabla_{W_{1}}L$ real number gradient of the loss function with respect to $W_{1}$
$\nabla_{b_{1}}L$ real number gradient of the loss function with respect to $b_{1}$
$\nabla_{W_{2}}L$ real number gradient of the loss function with respect to $W_{2}$
$\nabla_{b_{2}}L$ real number gradient of the loss function with respect to $b_{2}$
$\eta$ Positive real number learning rate
Table A3: Definition of Functions
@@
ReLU(x)(x) 1 rectified linear unit (ReLU) activation function
Identity(x)(x) 1 identity function
maxmax multivariate maximum function
[…]
ReLU’$(x)$ 1 derivative of ReLU(x)(x)
Target(X)(X) 1 objective function for input vector $X$
Table A4: Definition of Predicates
@@
[…]
A.3 Formal Representations of Information at Each Stage of Model M
@@ A.3.1 Initial information of M
$S_{oo}(oo,To_{h})=\{\varphi o_{1},\varphi o_{2},\varphi o_{3},\varphi o_{4},\varphi o_{5},\varphi o_{6},\varphi o_{7},\varphi o_{8},\varphi o_{9},\varphi o_{10},\varphi o_{11}\}$ (A.14)
@@
    $\varphi o_{1}=$ $\text{NetworkStructure}^{2}{(layers,3)}\land\text{LayerSize}^{2}(1,2)\land$ (A.15)
    $\text{LayerSize}^{2}(2,3)\land\text{LayerSize}^{2}(3,2)$
@@
    $\varphi o_{2}$ $=\forall i\forall j(BelongsTo^{2}(i,\{1,2,3\})\land BelongsTo^{2}(j,\{1,2\})\rightarrow\exists v(Eq^{2}(w_{ij}^{(o)1},v)))$ (A.16)
    $\varphi o_{3}$ $=\forall i\forall j(BelongsTo^{2}(i,\{1,2\})\land BelongsTo^{2}(j,\{1,2,3\})\rightarrow\exists v(Eq^{2}(w_{ij}^{(o)2},v)))$ (A.17)
@@
    $\varphi o_{4}=\forall iBelongsTo^{2}(i,\{1,2,3\})\rightarrow\exists v(Eq^{2}(b_{i}^{(o)1},v))$ (A.18)
    $\varphi o_{5}=\forall iBelongsTo^{2}(i,\{1,2\})\rightarrow\exists v(Eq^{2}(b_{i}^{(o)2},v))$ (A.19)
@@
    $\varphi o_{6}=\forall x(Eq^{2}(\text{ReLU}(x),max(0,x)))\land\forall x(Eq^{2}(\text{Identity}(x),x))$ (A.20)
@@
    $\varphi o_{7}=$ $\forall X\forall t\forall H(InputAt^{2}(X,t)\land BelongsTo^{2}(X,\mathbb{R}^{2})\rightarrow$
    $\exists H(BelongsTo^{2}(H,\mathbb{R}^{3})\land HiddenAt^{2}(H,t+\delta)\land$
    $\forall i(BelongsTo^{2}(i,\{1,2,3\})\rightarrow Eq^{2}(h_{i},\text{ReLU}\left(\sum_{j=1}^{2}w_{ij}^{(o)1}\times x_{j}+b_{i}^{(o)1}\right))))$ (A.21)
    $\varphi o_{8}=$ $\forall H\forall t(HiddenAt^{2}(H,t+\delta)\land BelongsTo^{2}(H,\mathbb{R}^{3})\rightarrow$
    $\exists Y(BelongsTo^{2}(Y,\mathbb{R}^{2})\land(OutputAt^{2}(Y,t+2\delta)\land$
    $\forall i(BelongsTo^{2}(i,\{1,2\})\rightarrow Eq^{2}(y_{i},\text{Identity}\left(\sum_{j=1}^{3}w_{ij}^{(o)2}\times h_{j}+b_{i}^{(o)2}\right)))))$ (A.22)
@@
    $\varphi o_{9}$ $=\forall X\forall Y_{t}\forall t(InputAt^{3}(X,Y_{t},t)\land BelongsTo^{2}(Y_{t},\mathbb{R}^{2})\land\varphi o_{7}\land\varphi o_{8}\land$
    $Eq^{2}(L,\text{MSE}(Y,Y_{t}))\land Eq^{2}(\text{MSE}(Y,Y_{t}),\frac{1}{2}\sum_{i=1}^{2}(y_{i}-y_{ti})^{2}))$ (A.23)
    $\varphi o_{10}=$ $\varphi o_{7}\land\varphi o_{8}\land\varphi o_{9}\land\text{Differentiable}^{2}(\text{MSE}(Y,Y_{t}),W_{1}^{(o)})\land$
    $\text{Differentiable}^{2}(\text{MSE}(Y,Y_{t}),b_{1}^{(o)})\land\text{Differentiable}^{2}(\text{MSE}(Y,Y_{t}),W_{2}^{(o)})\land$
    $\text{Differentiable}^{2}(\text{MSE}(Y,Y_{t}),b_{2}^{(o)})\land$
    $Eq^{2}(\nabla_{W^{(o)}_{1}}L,((W_{2}^{(o)})^{T}(Y-Y_{t})\odot\text{ReLU}^{\prime}(W_{1}^{(o)}X+b_{1}^{(o)}))X^{T})\land$
    $Eq^{2}(\nabla_{b^{(o)}_{1}}L,((W_{2}^{(o)})^{T}(Y-Y_{t})\odot\text{ReLU}^{\prime}(W_{1}^{(o)}X+b_{1}^{(o)}))\land$
    $Eq^{2}(\nabla_{W^{(o)}_{2}}L,(Y-Y_{t})(\text{ReLU}(W_{1}^{(o)}X+b_{1}^{(o)}))^{T})\land Eq^{2}(\nabla_{b^{(o)}_{2}}L,(Y-Y_{t}))$ (A.24)
@@
    $\varphi o_{11}=$ $\varphi o_{7}\land\varphi o_{8}\land\varphi o_{9}\land\varphi o_{10}\land Eq^{2}(W_{1}^{(u)},W_{1}^{(o)}-\eta\nabla_{W^{(o)}_{1}}L)\land$
    $Eq^{2}(b_{1}^{(u)},b_{1}^{(o)}-\eta\nabla_{b^{(o)}_{1}}L)\land Eq^{2}(W_{2}^{(u)},W_{2}^{(o)}-\eta\nabla_{W^{(o)}_{2}}L)\land$
    $Eq^{2}(b_{2}^{(u)},b_{2}^{(o)}-\eta\nabla_{b^{(o)}_{2}}L)$ (A.25)
@@
$\sup(To_{h})\leq\inf(To_{m})$ (A.26)
The state $S_{M}(M,To_{m})$ is a one-to-one enabling mapping of the state $S_{oo}(oo,To_{h})$ onto M, that is,
$S_{M}(M,To_{m})=\{\psi o_{1},\psi o_{2},\psi o_{3},\psi o_{4},\psi o_{5},\psi o_{6},\psi o_{7},\psi o_{8},\psi o_{9},\psi o_{10},\psi o_{11}\}$ (A.27)
@@
$Io(\varphi o_{i})\underline{\triangleright}\psi o_{i},(1\leq i\leq 11)$ (A.28)
Here, the information enables the mapping from the initial ontological state of model M to its carrier state, that is, the initial state of M. In practice, this mapping is a complex process involving hardware implementation, software development, and hardware-software integration, inevitably requiring consumption of time and energy resources, and subject to various sources of noise. However, to simplify the discussion, we temporarily disregard these factors and explain the initial state of M using a completely ideal, noise-free, and recoverable information model. The subsequent discussions on other types of information will follow this approach and will not be repeated.
A.3.2 Training Information of M
@@
$S_{ot}(ot,Tt_{h})=\{\varphi t_{1},\varphi t_{2}\}$ (A.30)
@@
    $\varphi t_{1}=\forall X\forall Y_{t}(BelongsTo^{2}(X,\mathbb{R}^{2})\land BelongsTo^{2}(Y_{t},\mathbb{R}^{2}))$ (A.31)
@@
    $\varphi t_{2}=\forall X\forall Y_{t}Eq^{2}(Target(X),Y_{t})$ (A.32)
@@
$It(\varphi t_{i})\underline{\triangleright}\psi t_{i}(1\leq i\leq 2)$ (A.33)
A.3.3 Optimized Information After M Has Learned from Training
Based on MLT-MF, after learning the training information $It$, model M updates its initial information $I_{0}$ to a new, optimized information state:
$I_{u}=(ou,Tu_{h},S_{ou},M,Tu_{m},S_{M})$ (A.34)
@@ A.3.3 Optimized Information After M Has Learned from Training
$S_{ou}(ou,Tu_{h})=\{\varphi u_{1},\varphi u_{6},\varphi u_{7},\varphi u_{8},\varphi u_{9},\varphi u_{10},\varphi u_{11}\}$ (A.35)
Since the structure and algorithms of M remain unchanged before and after training, $\varphi u_{1}$, $\varphi u_{6}$, and $\varphi u_{7}$ are exactly the same as $\varphi o_{1}$, $\varphi o_{6}$, and $\varphi o_{9}$ respectively. Moreover, as the weights and biases after training are already assigned values in the original information (A.25), there is no need to include equivalent formulas for $\varphi o_{2}$, $\varphi o_{3}$, $\varphi o_{4}$, $\varphi o_{5}$ in $S_{ou}(ou,Tu_{h})$. For the remaining formulas $\varphi u_{7}$, $\varphi u_{8}$, $\varphi u_{9}$, $\varphi u_{10}$, and $\varphi u_{11}$, it suffices to systematically replace $W_{k}^{(o)}$, $w_{ij}^{(o)k}$, $b_{k}^{(o)}$, $b_{i}^{(o)k}$, $W_{k}^{(u)}$, $w_{ij}^{(u)k}$, $b_{k}^{(u)}$ and $b_{i}^{(u)k}$ in $\varphi_{7}$, $\varphi_{8}$, $\varphi_{10}$, and $\varphi_{11}$ with the corresponding symbols $W_{k}^{(u)}$, $w_{ij}^{(u)k}$, $b_{k}^{(u)}$, $b_{i}^{(u)k}$, $W_{k}^{(v)}$, $w_{ij}^{(v)k}$, $b_{k}^{(v)}$ and $b_{i}^{(v)k}$ in $S_{ou}(ou,Tu_{h})$.

### v10

@@ C.1 A Simple Feedforward Neural Network Model
$W_{1}=\begin{bmatrix}w_{11}^{1}&w_{12}^{1}\\ w_{21}^{1}&w_{22}^{1}\\ w_{31}^{1}&w_{32}^{1}\end{bmatrix},\quad b_{1}=\begin{bmatrix}b_{1}^{1}\\ b_{2}^{1}\\ b_{3}^{1}\end{bmatrix}$ (C.3)
$W_{2}=\begin{bmatrix}w_{11}^{2}&w_{12}^{2}&w_{13}^{2}\\ w_{21}^{2}&w_{22}^{2}&w_{23}^{2}\end{bmatrix},\quad b_{2}=\begin{bmatrix}b_{1}^{2}\\ b_{2}^{2}\end{bmatrix}$ (C.4)
@@
$X=\begin{bmatrix}x_{1}\\ x_{2}\end{bmatrix},\quad Y=\begin{bmatrix}y_{1}\\ y_{2}\end{bmatrix},\quad Y_{t}=\begin{bmatrix}y_{t1}\\ y_{t2}\end{bmatrix}\text{ and }Y_{t}=Target(X)$ (C.5)
$L=MSE\left(Y,Y_{t}\right)=\frac{1}{2}\left\|Y-Y_{t}\right\|^{2}=\frac{1}{2}(\left(y_{1}-y_{t1}\right)^{2}+\left(y_{2}-y_{t2}\right)^{2})$ (C.6)
@@
$\nabla_{W_{2}}L=(Y-Y_{t}){(ReLU(W_{1}X+b_{1}))}^{T}$ (C.7)
$\nabla_{b_{2}}L=(Y-Y_{t})$ (C.8)
$\nabla_{W_{1}}L=({W_{2}}^{T}(Y-Y_{t})\bigodot{ReLU}^{\prime}(W_{1}X+b_{1}))X^{T}$ (C.9)
$\nabla_{b_{1}}L={W_{2}}^{T}(Y-Y_{t})\bigodot{ReLU}^{\prime}(W_{1}X+b_{1})$ (C.10)
[…]
$W_{1}^{\text{new}}=W_{1}-\eta\nabla_{W_{1}}L,\quad b_{1}^{\text{new}}=b_{1}-\eta\nabla_{b_{1}}L$ (C.11)
$W_{2}^{\text{new}}=W_{2}-\eta\nabla_{W_{2}}L,\quad b_{2}^{\text{new}}=b_{2}-\eta\nabla_{b_{2}}L$ (C.12)
@@
C.2 Symbol Definitions
MLT-MF uses a formal system to express the operational logic of each phase throughout the model’s full lifecycle. Defining the symbols is a prerequisite for formal modeling. To this end, Tables C.1–C.4 provide definitions for the constants, variables, functions, and predicates used in the formal representation of model M.
Table 2: Definition of Constants
@@ C.2 Symbol Definitions
Table 3: Definition of Variables - Part I: Logical Formulas and Basic Variables
@@
$\varphi o_{i}$ WFFs well-formed formulas in initial model ontological state
$\psi o_{i}$ WFFs well-formed formulas in initial model carrier state
$\varphi t_{i}$ WFFs well-formed formulas in training ontological state
$\psi t_{i}$ WFFs well-formed formulas in training carrier state
$\varphi u_{i}$ WFFs well-formed formulas in optimized model ontological state
$\psi u_{i}$ WFFs well-formed formulas in optimized model carrier state
$\varphi q_{i}$ WFFs well-formed formulas in input ontological state
$\psi q_{i}$ WFFs well-formed formulas in input carrier state
$\varphi r_{i}$ WFFs well-formed formulas in output ontological state
$\psi r_{i}$ WFFs well-formed formulas in output carrier state
$\varphi v_{i}$ WFFs well-formed formulas in re-optimized model ontological state
$\psi v_{i}$ WFFs well-formed formulas in re-optimized model carrier state
@@
Table 4: Definition of Variables - Part II: Neural Network Parameters
Symbol Type Definition
$W_{k}^{(o)}$ matrix weight matrix from k-th to (k+1)-th layer (initial)
$w_{ij}^{(o)k}$ real number weight from j-th to i-th neuron (initial)
$b_{k}^{(o)}$ vector bias vector of k-th layer (initial)
$b_{i}^{(o)k}$ real number bias of i-th neuron in k-th layer (initial)
$W_{k}^{(u)}$ matrix weight matrix from k-th to (k+1)-th layer (optimized)
$w_{ij}^{(u)k}$ real number weight from j-th to i-th neuron (optimized)
$b_{k}^{(u)}$ vector bias vector of k-th layer (optimized)
$b_{i}^{(u)k}$ real number bias of i-th neuron in k-th layer (optimized)
$W_{k}^{(v)}$ matrix weight matrix from k-th to (k+1)-th layer (re-optimized)
$w_{ij}^{(v)k}$ real number weight from j-th to i-th neuron (re-optimized)
$b_{k}^{(v)}$ vector bias vector of k-th layer (re-optimized)
$b_{i}^{(v)k}$ real number bias of i-th neuron in k-th layer (re-optimized)
$W_{k}^{(w)}$ matrix weight matrix from k-th to (k+1)-th layer (further re-optimized)
$w_{ij}^{(w)k}$ real number weight from j-th to i-th neuron (further re-optimized)
$b_{k}^{(w)}$ vector bias vector of k-th layer (further re-optimized)
$b_{i}^{(w)k}$ real number bias of i-th neuron in k-th layer (further re-optimized)
Table 5: Definition of Variables - Part III: Vectors and Optimization Parameters
Symbol Type Definition
@@
$x_{i}$ real number i-th element of input vector
$h_{i}$ real number i-th element of hidden layer vector
$y_{i}$ real number i-th element of output vector
$y_{ti}$ real number i-th element of output reference vector
@@
$\nabla_{W_{1}}L$ real number gradient of loss function w.r.t. $W_{1}$
$\nabla_{b_{1}}L$ real number gradient of loss function w.r.t. $b_{1}$
$\nabla_{W_{2}}L$ real number gradient of loss function w.r.t. $W_{2}$
$\nabla_{b_{2}}L$ real number gradient of loss function w.r.t. $b_{2}$
$\eta$ positive real learning rate
Table 6: Definition of Functions
@@
$ReLU(x)$ 1 rectified linear unit (ReLU) activation function
$Identity(x)$ 1 identity function
$\max$ multivariate maximum function
$MSE(Y,Y_{t})$ 2 mean squared error between vector $Y$ and $Y_{t}$
${ReLU}^{\prime}(x)$ 1 derivative of $ReLU(x)$
$Target(X)$ 1 objective function for input vector $X$
Table 7: Definition of Predicates
@@
${NetworkStructure}^{2}(layers,k)$ model structure has k layers
${LayerSize}^{2}(k,j)$ the k-th layer has j neurons
${Eq}^{2}(x,y)$ $x$ equals $y$
${InputAt}^{2}(X,t)$ input $X$ at time $t$
${HiddenAt}^{2}(H,t)$ hidden layer value is $H$ at time $t$
${OutputAt}^{2}(Y,t)$ output $Y$ at time $t$
${InputAt}^{3}(X,Y_{t},t)$ input vector $X$ and label vector $Y_{t}$ at time $t$
${Differentiable}^{2}(f,v)$ function $f$ is differentiable with respect to $v$
${BelongsTo}^{2}(x,X)$ element $x$ belongs to set $X$
C.3 Formal Representations of Information at Each Stage of Model M
@@
$S_{oo}\left(oo,{To}_{h}\right)=\left\{\varphi o_{1},\varphi o_{2},\varphi o_{3},\varphi o_{4},\varphi o_{5},\varphi o_{6},\varphi o_{7},\varphi o_{8},\varphi o_{9},\varphi o_{10},\varphi o_{11}\right\}$ (C.14)
@@
$\varphi o_{1}={NetworkStructure}^{2}(layers,3)\land{LayerSize}^{2}(1,2)\land{LayerSize}^{2}(2,3)\land{LayerSize}^{2}(3,2)$ (C.15)
@@
$\varphi o_{2}=\forall i{\forall j(BelongsTo}^{2}\left(i,\left\{1,2,3\right\}\right){\land BelongsTo}^{2}\left(j,\left\{1,2\right\}\right)\rightarrow{\exists v(Eq}^{2}\left(w_{ij}^{(o)1},v\right))$ (C.16)
$\varphi o_{3}=\forall i{\forall j(BelongsTo}^{2}\left(i,\left\{1,2\right\}\right){\land BelongsTo}^{2}\left(j,\left\{1,2,3\right\}\right)\rightarrow{\exists v(Eq}^{2}\left(w_{ij}^{(o)2},v\right))$ (C.17)
@@
$\varphi o_{4}=\forall i{BelongsTo}^{2}\left(i,\left\{1,2,3\right\}\right)\rightarrow{\exists v(Eq}^{2}\left(b_{i}^{(o)1},v\right))$ (C.18)
$\varphi o_{5}=\forall i{BelongsTo}^{2}\left(i,\left\{1,2\right\}\right)\rightarrow{\exists v(Eq}^{2}\left(b_{i}^{(o)2},v\right))$ (C.19)
@@
$\varphi o_{6}=\forall x({Eq}^{2}(ReLU(x),max(0,x)))\land\forall x({Eq}^{2}(Identity(x),x))$ (C.20)
@@
$\begin{split}\varphi o_{7}=&\forall X\forall t({InputAt}^{2}(X,t)\land{BelongsTo}^{2}\left(X,\mathbb{R}^{2}\right)\rightarrow\exists H({BelongsTo}^{2}\left(H,\mathbb{R}^{3}\right)\\ &\land{HiddenAt}^{2}(H,t+\delta)\land\forall i({BelongsTo}^{2}\left(i,\left\{1,2,3\right\}\right)\\ &\rightarrow{Eq}^{2}(h_{i},ReLU\left(\sum_{j=1}^{2}w_{ij}^{(o)1}\times x_{j}+b_{i}^{(o)1}\right)))))\end{split}$ (C.21)
$\begin{split}\varphi o_{8}=&\forall H\forall t({HiddenAt}^{2}(H,t+\delta)\land{BelongsTo}^{2}\left(H,\mathbb{R}^{3}\right)\rightarrow\exists Y({BelongsTo}^{2}\left(Y,\mathbb{R}^{2}\right)\\ &\land{OutputAt}^{2}(Y,t+2\delta)\land\forall i({BelongsTo}^{2}\left(i,\left\{1,2\right\}\right)\\ &\rightarrow{Eq}^{2}\left(y_{i},Identity\left(\sum_{j=1}^{3}w_{ij}^{(o)2}\times h_{j}+b_{i}^{(o)2}\right)\right))))\end{split}$ (C.22)
@@
$\begin{split}\varphi o_{9}=&\forall X\forall Y_{t}\forall t({InputAt}^{3}\left(X,Y_{t},t\right)\land{BelongsTo}^{2}\left(Y_{t},\mathbb{R}^{2}\right)\land\varphi o_{7}\land\varphi o_{8}\\ &{\land Eq}^{2}\left(L,MSE\left(Y,Y_{t}\right)\right)\land{Eq}^{2}\left(MSE\left(Y,Y_{t}\right),\frac{1}{2}\sum_{i=1}^{2}\left(y_{i}-y_{ti}\right)^{2}\right))\end{split}$ (C.23)
$\varphi o_{10}$ $=\varphi o_{7}\land\varphi o_{8}\land\varphi o_{9}$
$\quad\land{Differentiable}^{2}(MSE(Y,Y_{t}),W_{1}^{(o)})$
$\quad\land{Differentiable}^{2}(MSE(Y,Y_{t}),b_{1}^{(o)})$
$\quad\land{Differentiable}^{2}(MSE(Y,Y_{t}),W_{2}^{(o)})$
$\quad\land{Differentiable}^{2}(MSE(Y,Y_{t}),b_{2}^{(o)})$
$\quad\land{Eq}^{2}(\nabla_{W_{1}^{(o)}}L,$
$\qquad({(W_{2}^{(o)})}^{T}(Y-Y_{t})\bigodot{ReLU}^{\prime}(W_{1}^{(o)}X+b_{1}^{(o)}))X^{T})$
$\quad\land{Eq}^{2}(\nabla_{b_{1}^{(o)}}L,$
$\qquad{(W_{2}^{(o)})}^{T}(Y-Y_{t})\bigodot{ReLU}^{\prime}(W_{1}^{(o)}X+b_{1}^{(o)}))$
$\quad\land{Eq}^{2}(\nabla_{W_{2}^{(o)}}L,(Y-Y_{t}){(ReLU(W_{1}^{(o)}X+b_{1}^{(o)}))}^{T})$
$\quad\land{Eq}^{2}(\nabla_{b_{2}^{(o)}}L,(Y-Y_{t}))$ (C.24)
@@
$\begin{split}\varphi o_{11}=&\varphi o_{7}\land\varphi o_{8}\land\varphi o_{9}\land\varphi o_{10}\land{Eq}^{2}(W_{1}^{(u)},W_{1}^{(o)}-\eta\nabla_{W_{1}^{(o)}}L)\\ &\land{Eq}^{2}(b_{1}^{(u)},b_{1}^{(o)}-\eta\nabla_{b_{1}^{(o)}}L)\land{Eq}^{2}(W_{2}^{(u)},W_{2}^{(o)}-\eta\nabla_{W_{2}^{(o)}}L)\\ &\land{Eq}^{2}(b_{2}^{(u)},b_{2}^{(o)}-\eta\nabla_{b_{2}^{(o)}}L)\end{split}$ (C.25)
@@
$sup({To}_{h})\leq inf({To}_{m})$ (C.26)
The state $S_{M}(M,{To}_{m})$ is a one-to-one enabling mapping of the state $S_{oo}\left(oo,{To}_{h}\right)$ onto M, that is,
$S_{M}\left(M,{To}_{m}\right)=\left\{\psi o_{1},\psi o_{2},\psi o_{3},\psi o_{4},\psi o_{5},\psi o_{6},\psi o_{7},\psi o_{8},\psi o_{9},\psi o_{10},\psi o_{11}\right\}$ (C.27)
@@
$Io\left(\varphi o_{i}\right)\trianglerighteq\psi o_{i}(1\leq i\leq 11)$ (C.28)
Here, the information $Io$ enables the mapping from the initial ontological state of model M to its carrier state, that is, the initial state of M. In practice, this mapping is a complex process involving hardware implementation, software development, and hardware-software integration, inevitably requiring consumption of time and energy resources, and subject to various sources of noise. However, to simplify the discussion, we temporarily disregard these factors and explain the initial state of M using a completely ideal, noise-free, and recoverable information model. The subsequent discussions on other types of information will follow this approach and will not be repeated.
(2) Training Information of M
@@
$S_{ot}\left(ot,{Tt}_{h}\right)=\left\{\varphi t_{1},\varphi t_{2}\right\}$ (C.30)
@@
$\varphi t_{1}=\forall X\forall Y_{t}({BelongsTo}^{2}(X,\mathbb{R}^{2})\land{BelongsTo}^{2}(Y_{t},\mathbb{R}^{2}))$ (C.31)
@@
$\varphi t_{2}=\forall X\forall Y_{t}{Eq}^{2}(Target(X),Y_{t})$ (C.32)
@@
$It\left(\varphi t_{i}\right)\trianglerighteq\psi t_{i}(1\leq i\leq 2)$ (C.33)
(3) Optimized Information After M Has Learned from Training
Based on MLT-MF, after learning the training information $It$, model M updates its initial information $Io$ to a new, optimized information state:
$Iu=\left\langle ou,{Tu}_{h},S_{ou},M,{Tu}_{m},S_{M}\right\rangle$ (C.34)
@@
$S_{ou}\left(ou,{Tu}_{h}\right)=\left\{\varphi u_{1},\varphi u_{6},\varphi u_{7},\varphi u_{8},\varphi u_{9},\varphi u_{10},\varphi u_{11}\right\}$ (C.35)
Since the structure and algorithms of M remain unchanged before and after training, $\varphi u_{1},\varphi u_{6}$, and $\varphi u_{9}$ are exactly the same as $\varphi o_{1},\varphi o_{6}$, and $\varphi o_{9}$ respectively. Moreover, as the weights and biases after training are already assigned values in the original information (A.25), there is no need to include equivalent formulas for $\varphi o_{2},\varphi o_{3},\varphi o_{4}$, or $\varphi o_{5}$ in $S_{ou}\left(ou,{Tu}_{h}\right)$. For the remaining formulas $\varphi u_{7},\varphi u_{8},\varphi u_{10}$, and $\varphi u_{11}$, it suffices to systematically replace $W_{k}^{(o)}$, $w_{ij}^{(o)k}$, $b_{k}^{(o)}$, $b_{i}^{(o)k}$, $W_{k}^{(u)}$, $w_{ij}^{(u)k}$, $b_{k}^{(u)}$ and $b_{i}^{(u)k}$ in $\varphi o_{7},\varphi o_{8},\varphi o_{10}$, and $\varphi o_{11}$ with the corresponding symbols $W_{k}^{(u)}$, $w_{ij}^{(u)k}$, $b_{k}^{(u)}$, $b_{i}^{(u)k}$, $W_{k}^{(v)}$, $w_{ij}^{(v)k}$, $b_{k}^{(v)}$ and $b_{i}^{(v)k}$ in $S_{ou}\left(ou,{Tu}_{h}\right)$.